# TransXNet: Learning Both Global and Local Dynamics with a Dual Dynamic Token Mixer for Visual Recognition

### v3

@@ Abstract
Recent studies have integrated convolutions into transformers to introduce inductive bias and improve generalization performance. However, the static nature of conventional convolution prevents it from dynamically adapting to input variations, resulting in a representation discrepancy between convolution and self-attention as self-attention calculates attention matrices dynamically. Furthermore, when stacking token mixers that consist of convolution and self-attention to form a deep network, the static nature of convolution hinders the fusion of features previously generated by self-attention into convolution kernels. These two limitations result in a sub-optimal representation capacity of the constructed networks. To find a solution, we propose a lightweight Dual Dynamic Token Mixer (D-Mixer) to simultaneously learn global and local dynamics, that is, mechanisms that compute weights for aggregating global contexts and local details in an input-dependent manner. D-Mixer works by applying an efficient global attention module and an input-dependent depthwise convolution separately on evenly split feature segments, endowing the network with strong inductive bias and an enlarged effective receptive field. We use D-Mixer as the basic building block to design TransXNet, a novel hybrid CNN-Transformer vision backbone network that delivers compelling performance. In the ImageNet-1K image classification task, TransXNet-T surpasses Swin-T by 0.3% in top-1 accuracy while requiring less than half of the computational cost. Furthermore, TransXNet-S and TransXNet-B exhibit excellent model scalability, achieving top-1 accuracy of 83.8% and 84.6% respectively, with reasonable computational costs. Additionally, our proposed network architecture demonstrates strong generalization capabilities in various dense prediction tasks, outperforming other state-of-the-art networks while having lower computational costs. Code is publicly available at https://github.com/LMMMEng/TransXNet.
@@ I Introduction
On the basis of the above discussions, we introduce a novel Dual Dynamic Token Mixer (D-Mixer) to learn both global and local dynamics, namely, mechanisms that compute weights for aggregating global and local features in an input-dependent way. Specifically, the input features are split into two half segments, which are respectively processed by an Overlapping Spatial Reduction Attention module and an Input-dependent Depthwise Convolution. The resulting two outputs are then concatenated together. Such a simple design can make a network see global contextual information while injecting effective inductive bias. As shown in Fig. 1, our method stands out among its competitors, yielding the largest ERF. In addition, zoom-in views (last column) reveal that our proposed mixer has remarkable local sensitivity in addition to non-local attention. We further introduce a Multi-scale Feed-forward Network (MS-FFN) that explores multi-scale information during token aggregation. By hierarchically stacking basic blocks composed of a D-Mixer and an MS-FFN, we construct a versatile backbone network called TransXNet for visual recognition. As illustrated in Fig. 2, our method showcases superior performance when compared to recent state-of-the-art (SOTA) methods in ImageNet-1K [14] image classification. In particular, our TransXNet-T achieves 81.6% top-1 accuracy with only 1.8G FLOPs and 12.8M Parameters (Params), outperforming Swin-T while incurring less than half of its computational cost. Also, our TransXNet-S/B models achieve 83.8%/84.6% top-1 accuracy, surpassing recently proposed InternImage [15] while incurring less computational cost.
@@ II-D Dynamic Weights
Dynamic weight is a powerful factor for the superiority of self-attention, enabling it to extract features dynamically according to the input, in addition to its long-range modeling capability. Similarly, dynamic convolution has been shown to be effective in improving the performance of CNN models [37, 38, 39, 40] by extracting more discriminative local features with input-dependent filters. Among these methods, Han et al. have [40] demonstrated that replacing the shifted window attention modules in Swin Transformer with dynamic depthwise convolutions achieves better results with lower computational cost.
@@ III-B1 Input-dependent Depthwise Convolution
To inject inductive bias and perform local feature aggregation in a dynamic input-dependent way, we propose a new type of dynamic depthwise convolution, termed Input-dependent Depthwise Convolution (IDConv). As shown in Fig. 4 (b), taking an input feature map $\mathbf{X}\in\mathbb{R}^{C\times H\times W}$, adaptive average pooling is used to aggregate spatial contexts, compressing spatial dimension to $\mathit{K}^{2}$, which is then forwarded into two sequential 1$\times$1 convolutions, yielding attention maps $\mathbf{A}^{\prime}\in\mathbb{R}^{(G\times C)\times{K}^{2}}$, where $G$ denotes the number of attention groups. Then, $\mathbf{A}^{\prime}$ is reshaped into $\mathbb{R}^{G\times C\times{K}^{2}}$ and a softmax function is employed over the ${G}$ dimension, thus generating attention weights $\mathbf{A}\in\mathbb{R}^{G\times C\times{K}^{2}}$. Finally, $\mathbf{A}$ is element-wise multiplied with a set of learnable parameters $\mathbf{P}\in\mathbb{R}^{G\times C\times K^{2}}$, and the output is summed over the ${G}$ dimension, resulting in input-dependent depthwise convolution kernels $\mathbf{W}\in\mathbb{R}^{C\times K^{2}}$, which can be expressed as:
@@ V Limitations
Our ablation study suggests that a fixed 1:1 ratio between the numbers of channels allocated to self-attention and dynamic convolutions in all stages yields a favorable trade-off. However, we speculate that employing different ratios at different stages may further improve performance and reduce computational cost. Regarding the model design, our TransXNet series are manually stacked, and there exist potentially inefficient operators in the building blocks (e.g., multi-scale depthwise convolutions in MS-FFN). As a result, our model exhibits limited advantages in terms of speed when compared to other models with similar GFLOPs. Nonetheless, these inefficiencies can be mitigated through techniques such as Neural Architecture Search (NAS) [63] and specialized implementation engineering, which we plan to explore in our future work.

### v4

@@ Abstract
Recent studies have integrated convolutions into transformers to introduce inductive bias and improve generalization performance. However, the static nature of conventional convolution prevents it from dynamically adapting to input variations, resulting in a representation discrepancy between convolution and self-attention as the latter computes attention maps dynamically. Furthermore, when stacking token mixers that consist of convolution and self-attention to form a deep network, the static nature of convolution hinders the fusion of features previously generated by self-attention into convolution kernels. These two limitations result in a sub-optimal representation capacity of the entire network. To find a solution, we propose a lightweight Dual Dynamic Token Mixer (D-Mixer) to simultaneously learn global and local dynamics via computing input-dependent global and local aggregation weights. D-Mixer works by applying an efficient global attention module and an input-dependent depthwise convolution separately on evenly split feature segments, endowing the network with strong inductive bias and an enlarged receptive field. We use D-Mixer as the basic building block to design TransXNet, a novel hybrid CNN-Transformer vision backbone network that delivers compelling performance. In the ImageNet-1K classification, TransXNet-T surpasses Swin-T by 0.3% in top-1 accuracy while requiring less than half of the computational cost. Furthermore, TransXNet-S and TransXNet-B exhibit excellent model scalability, achieving top-1 accuracy of 83.8% and 84.6% respectively, with reasonable computational costs. Additionally, our proposed network architecture demonstrates strong generalization capabilities in various dense prediction tasks, outperforming other state-of-the-art networks while having lower computational costs. Code is publicly available at https://github.com/LMMMEng/TransXNet.
@@ I Introduction
On the basis of the above discussions, we introduce a novel Dual Dynamic Token Mixer (D-Mixer) to learn both global and local dynamics, namely, mechanisms that compute weights for aggregating global and local features in an input-dependent way. Specifically, the input features are split into two half segments, which are respectively processed by an Overlapping Spatial Reduction Attention module and an Input-dependent Depthwise Convolution. The resulting two outputs are then concatenated together. Such a simple design can make a network see global contextual information while injecting effective inductive bias. As shown in Fig. 2, our method stands out among its competitors, yielding the largest ERF. In addition, zoom-in views (last column) reveal that our proposed mixer has remarkable local sensitivity in addition to non-local attention. We further introduce a Multi-scale Feed-forward Network (MS-FFN) that explores multi-scale information during token aggregation. By hierarchically stacking basic blocks composed of a D-Mixer and an MS-FFN, we construct a versatile backbone network called TransXNet for visual recognition. As illustrated in Fig. 1, our method showcases superior performance when compared to recent state-of-the-art (SOTA) methods in ImageNet-1K [14] image classification. In particular, our TransXNet-T achieves 81.6% top-1 accuracy with only 1.8 GFLOPs and 12.8M Parameters (Params), outperforming Swin-T while incurring less than half of its computational cost. Also, our TransXNet-S/B models achieve 83.8%/84.6% top-1 accuracy, surpassing the strong InternImage [15] while incurring less computational cost.
@@ II-D Dynamic Weights
Dynamic weight is a powerful factor for the superiority of self-attention, enabling it to extract features dynamically according to the input, in addition to its long-range modeling capability. Similarly, dynamic convolution has been shown to be effective in improving the performance of CNN models [37, 38, 39, 40] by extracting more discriminative local features with input-dependent filters. Among these methods, Han et al. [40] have demonstrated that replacing the shifted window attention modules in Swin Transformer with dynamic depthwise convolutions achieves better results with lower computational cost.
@@ III-B1 Input-dependent Depthwise Convolution
To inject inductive bias and perform local feature aggregation in a dynamic input-dependent way, we propose a new type of dynamic depthwise convolution, termed Input-dependent Depthwise Convolution (IDConv). As shown in Fig. 4 (b), taking an input feature map $\mathbf{X}\in\mathbb{R}^{C\times H\times W}$, an adaptive average pooling layer is used to aggregate spatial contexts, compressing spatial dimension to $\mathit{K}^{2}$, which is then forwarded into two sequential 1$\times$1 convolutions, yielding attention maps $\mathbf{A}^{\prime}\in\mathbb{R}^{(G\times C)\times{K}^{2}}$, where $G$ denotes the number of attention groups. Then, $\mathbf{A}^{\prime}$ is reshaped into $\mathbb{R}^{G\times C\times{K}^{2}}$ and a softmax function is employed over the ${G}$ dimension, thus generating attention weights $\mathbf{A}\in\mathbb{R}^{G\times C\times{K}^{2}}$. Finally, $\mathbf{A}$ is element-wise multiplied with a set of learnable parameters $\mathbf{P}\in\mathbb{R}^{G\times C\times K^{2}}$, and the output is summed over the ${G}$ dimension, resulting in input-dependent depthwise convolution kernels $\mathbf{W}\in\mathbb{R}^{C\times K^{2}}$, which can be expressed as:
@@ V Limitations
Our ablation study suggests that a fixed 1:1 ratio between the numbers of channels allocated to self-attention and dynamic convolutions in all stages yields a favorable trade-off. However, we speculate that employing different ratios at different stages may further improve performance and reduce computational cost. Regarding the model design, our TransXNet series are manually stacked, and there exist potentially inefficient operators in the building blocks (e.g., multi-kernel depthwise convolutions in MS-FFN). As a result, our model exhibits limited advantages in terms of speed when compared to other models with similar GFLOPs. Nonetheless, these inefficiencies can be mitigated through techniques such as Neural Architecture Search (NAS) [63] and specialized implementation engineering, which we plan to explore in our future work.